# SegIC: Unleashing the Emergent Correspondence for In-Context Segmentation

### v1

@@ Abstract
In-context segmentation aims at segmenting novel images using a few labeled example images, termed as “in-context examples”, exploring content similarities between examples and the target. The resulting models can be generalized seamlessly to novel segmentation tasks, significantly reducing the labeling and training costs compared with conventional pipelines. However, in-context segmentation is more challenging than classic ones due to its meta-learning nature, requiring the model to learn segmentation rules conditioned on a few samples, not just the segmentation. Unlike previous work with ad-hoc or non-end-to-end designs, we propose SegIC, an end-to-end segment-in-context framework built upon a single vision foundation model (VFM). In particular, SegIC leverages the emergent correspondence within VFM to capture dense relationships between target images and in-context samples. As such, information from in-context samples is then extracted into three types of instructions, i.e. geometric, visual, and meta instructions, serving as explicit conditions for the final mask prediction. SegIC is a straightforward yet effective approach that yields state-of-the-art performance on one-shot segmentation benchmarks. Notably, SegIC can be easily generalized to diverse tasks, including video object segmentation and open-vocabulary segmentation. Code will be available at https://github.com/MengLcool/SEGIC.
1 Introduction
@@
We design geometric, visual, and meta instructions that explicitly transfer knowledge from in-context samples to the target to facilitate in-context segmentation.
SegIC demonstrates state-of-the-art performance on COCO-20, FSS-1000 and recent LVIS-92. Furthermore, SegIC achieves competitive performance on novel tasks including video object segmentation and open-vocabulary segmentation, without ever seeing their training data.
@@ 2 Related Work
Vision foundation models. Recent years have witnessed great progress in large-scale vision pre-training [50, 8, 16, 6, 2, 17, 25], serving as the cornerstone for high-capacity foundation models. These pre-training approaches can be broadly categorized into two directions: vision-only pretraining [17, 16, 8, 6, 45, 2] and vision-language pre-training [50, 21, 30, 64, 52]. For vision-only pre-training, models aim to distinguish image/patch-level entities from different views [6, 8, 16] or reconstruct masked images [2, 16] from raw images. In vision-language pre-training, models strive to align cross-modal features into a unified visual-semantic space [50, 21, 64, 30], showcasing great open-set performance due to the transferable capabilities of language. Unlike these approaches that perform pre-training in an unsupervised or weakly supervised manner, SAM [25] is pre-trained on a huge amount of labeled segmentation data and can segment various entities with precise prompts about locations. In this paper, we conduct extensive experiments using three types of pre-trained models as backbones to explore their potential for in-context segmentation. Interestingly, we observe that models with higher zero-shot semantic and geometric correspondence performance are more likely to be effectively utilized in our SegIC framework for in-context segmentation.
Unified vision models. Unified vision models have recently drawn significant attention due to their generalization capabilities and flexibilities. Unlike previous specialized vision models designed for specific datasets [51, 36, 15, 7, 60], vision generalists are tailored to handle multiple datasets [68, 39, 13, 40] and a wide range of tasks [69, 70, 20] within a single yet unified model. Recently, many studies [69, 70, 20, 25, 58, 59] have focused on developing techniques that unify segmentation tasks. In a similar spirit, our work also builds upon a unified output space for segmentation tasks. However, our goal is different—we aim to perform in-context segmentation that allows a model to effectively segment novel entities conditioned on a few samples.
@@
In-context learning. For the first time, GPT-3 [4] introduces a new learning paradigm known as in-context learning, which unifies various NLP tasks as text completion or question-answering tasks using provided prompts and examples. This approach enables language models to handle various tasks, including novel ones, by leveraging task examples, without requiring re-training or fine-tuning. Recent studies [58, 59, 35] explore this mechanism in vision tasks. Painter [58] and SegGPT [59] aim to achieve in-context segmentation through in-painting. They build upon the Mask Image Modeling (MIM) framework [17] to concatenate images and predictions into a 22 mosaic and make predictions by recovering the masked areas. In their pipelines, the vision backbone serves as both an image encoder and a mask decoder, which incurs significant computational costs. Moreover, they struggle to effectively leverage pre-trained models due to input shifts, leading to increased convergence challenges. Other approaches [66, 35] attempt using in-context segmentation via prompting SAM [25]. They build upon cross-image correspondences between in-context examples and target images by additional pre-trained models to generate prompts for SAM. However, these methods employ a two-stage pipeline, introducing redundancy and repeated computations. Consequently, if the model encounters limitations in one stage, it negatively impacts the final performance. In this work, we build an end-to-end in-context segmentation framework, leveraging the emergent correspondence using a single vision foundation model.
@@ 3.2 In-context Instruction Extraction
Visual instructions indicate the visual clues of the target entity. We use the mask label from the reference image to extract visual salient clues, , from reference features:
By doing so, only relevant information in reference features as indicated in the mask, including low-level (texture, appearance, etc.) and high-level (semantics) clues, is used.
@@ 3.3 Mask Decoding
where indicate the projection layers for , respectively. are the projected in-context instructions.
@@ 4 Experiments
| Method | one-shot segmentation | video object segmentation | semantic seg | open-vocab seg | |||||
COCO-20 |
FSS-1000 |
LVIS-92 |
DAVIS-17 |
YVOS-18 |
COCO |
ADE20k |
PC-459 |
A-847 |
mean mIoU |
mIoU |
mean mIoU |
G |
mIoU |
mIoU |
mIoU |
mIoU |
| few-shot seg specialist | |||||||||
| HSNet [42] |
41.2/ 41.7 |
86.5 |
17.4 |
n/a | n/a | n/a | n/a | n/a | n/a |
| VAT [18] |
41.3/ 42.9 |
90.3 |
18.5 |
n/a | n/a | n/a | n/a | n/a | n/a |
| FPTrans [65] |
47.0/ 56.5 |
n/a | n/a | n/a | n/a | n/a | n/a |
| VOS specialist | |||||||||
| AGAME [22] | n/a | n/a | n/a |
70.0 |
66.0 |
n/a | n/a | n/a | n/a |
| SWEM [33] | n/a | n/a | n/a |
84.3 |
82.8 |
n/a | n/a | n/a | n/a |
| XMem [11] | n/a | n/a | n/a |
87.7 |
86.1 |
n/a | n/a | n/a | n/a |
| semantic seg specialist | |||||||||
| MaskFormer [9] (Swin-L) | n/a | n/a | n/a | n/a | n/a |
64.8 |
54.1 |
n/a | n/a |
| Mask2Former [10] (Swin-L) | n/a | n/a | n/a | n/a | n/a |
67.4 |
56.1 |
n/a | n/a |
| MaskDINO [28] (Swin-L) | n/a | n/a | n/a | n/a | n/a |
56.6 |
n/a | n/a | |
| segmentation generalist | |||||||||
| OneFormer [20] (Swin-L) | n/a | n/a | n/a | n/a | n/a |
67.4 |
57.7 |
| UNINEXT [63] (ViT-L) | n/a | n/a | n/a |
77.2 |
78.1 |
| X-decoder [69] (DaViT-L) | n/a | n/a | n/a | n/a | n/a |
67.5 |
58.1 |
29.6 |
9.2 |
| SEEM [70] (DaViT-L) |
58.9 |
50.0 |
67.6 |
| in-context generalist | |||||||||
| Painter [58] (MAE-L) |
33.1 |
61.7 |
10.5 |
34.6 |
24.1 |
49.9 |
| SegGPT [59] (MAE-L) |
56.1 |
85.6 |
18.6 |
75.6 |
74.7 |
*39.6 |
| PerSAM [66] (SAM-G) | 23.0 |
71.2 |
11.5 |
60.3 |
n/a | n/a | n/a | n/a |
| PerSAM-F [25] (SAM-G) | 23.5 |
75.6 |
12.3 |
71.9 |
n/a | n/a | n/a | n/a |
| Matcher [35] (DINO-v2-G+SAM-G) | 52.7 |
87.0 |
33.0 |
79.5 |
n/a | n/a | n/a | n/a |
| SegIC (DINO-v2-L) |
76.1 |
86.8 |
44.6 |
71.4 |
62.7 |
#72.9 |
#55.5 |
#23.9 |
#14.8 |
| SegIC (DINO-v2-G) |
74.5 |
88.4 |
47.8 |
73.7 |
65.4 |
#74.0 |
#59.0 |
#24.0 |
#15.4 |
@@ 4.1 Training Data
COCO [32] is a widely used instance/semantic segmentation with 83K training samples and 41K samples of 80 categories. We use the 2014 split version to be consistent with the COCO-20 one-shot semantic segmentation benchmark.
@@ 4.2 Training Details
We implement SegIC in PyTorch and use 8 V100 GPUs for most of our experiments. Thanks to the frozen backbone, our method is extremely memory-efficient (with less than 10G memory cost in most of our experiments). SegIC is a single unified model that is jointly trained on mixed datasets, while evaluated on various datasets, separately. We utilize DINO-v2 [45] as the default vision foundation model, with a ViT-B for all ablations, and ViT-L/G for the main experiments. We employ an AdamW [37] optimizer with a weight decay of We set an initial learning rate as and multiply 0.1 at the 10 epoch during training. Each dataset is sampled uniformly with 160K samples per epoch in the main experiments and 80K samples for the ablations. We perform dataset augmentations on target images and reference images respectively. We use large-scale jittering augmentation for semantic segmentation datasets, and normal data augmentations, including random resizing cropping, color jittering, and random horizontal flipping, for instance segmentation datasets. We random sample 1 mask per image for semantic segmentation datasets during training, while up to 10 masks for instance segmentation. The size of a single image is padded to 896896.
@@ 4.3 Main Results
For the main experiments, we compare SegIC with other specialist/generalist methods on several benchmarks. We use SegIC of DINO-v2 [45] of large and giant versions as the backbone for the main experiments.
One-shot semantic segmentation. Following SegGPT [59], we have evaluated SegIC in two one-shot semantic segmentation settings: in-domain using COCO-20[43] (the training set of COCO-20 is a subset of ours) and out-of-domain with FSS-1000 [31] (without seeing any training samples of FSS-1000). As depicted in Tab. 1, SegIC has achieved state-of-the-art performance in both of these settings. To ensure a fair comparison, we report both out-of-domain (the former) and in-domain (the latter) performance of specialist models for COCO-20, split by / in Tab. 1. Additionally, for FSS-1000, we highlight the performance trained on FSS-1000 in gray and zero-shot results in black. Notably, SegIC outperforms previous generalist models by a significant margin (more than 20 of mean mIoU) on COCO-20 and achieves competitive results that are very close to specialist models on FSS-1000, even without ever being trained on it. Furthermore, we conduct experiments on LVIS-92 [35], a more challenging one-shot benchmark built upon LVIS [14]. On LVIS-92, SegIC surpasses Matcher, the previous best model, by a large margin (from 33.0 to 47.8).
| Method |
F0 |
F1 |
F2 |
F3 |
mean |
| HSNet [42] |
37.2 |
44.1 |
42.4 |
41.3 |
41.2 |
| VAT [18] |
39.0 |
43.8 |
42.6 |
39.7 |
41.3 |
| FPTrans [65] |
44.4 |
48.9 |
50.6 |
44.0 |
47.0 |
| MSANet [19] |
47.8 |
57.4 |
48.7 |
50.5 |
51.1 |
| SegIC |
55.8 |
54.7 |
52.4 |
51.4 |
53.6 |
@@
Since we include the evaluation categories in the training process for COCO-20 in Tab. 1, for a rigorous comparison, we follow its training setting that trains and tests our model on 4 splits [43] separately, avoiding seeing categories for evaluation. As shown in Tab. 2, our method still achieves state-of-the-art performance on COCO-20. Furthermore, as shown in the last row of Tab. 2, by joint training on COCO-excluded datasets (including ADE20k, LVIS, and FSS-1000), there is a significant performance gain across all splits. This further demonstrates the effectiveness of our in-context generalization capabilities.
@@
Zero-shot video object segmentation. Video object segmentation (VOS) aims to segment specific objects in video frames. In this work, we focus on the semi-supervised VOS setting, where the masks that appeared first time are given as references. We evaluate SegIC on the validation set of two commonly used VOS datasets: DAVIS-17 [47] and YouTube-VOS-18 [62]. We export two metrics commonly used in VOS for evaluation: the score for DAVIS-17 and score for YouTube-VOS-18, with their official evaluation servers or toolkits. As shown in Tab. 1, when compared to VOS specialist models, SegIC achieves competitive performance on VOS benchmarks, even without seeing any training videos. Furthermore, in comparison to segmentation generalist models, SegIC surpasses Painter [58], SEEM [70], and PerSAM [66] by a significant margin, and competes favorably with recent generalist models [59, 35]. Additionally, the VOS task pipeline in SegIC is relatively simple. We do not use any test time augmentation tricks (TTA) used in [11]. Moreover, it does not involve dense feature interaction at the patch level, as in SegGPT, and does not require the use of a pre-trained SAM for segmentation.
| Encoder |
COCO-20 |
FSS-1000 |
DAVIS-17 |
| DINO-v2-B [45] |
74.6 |
87.4 |
70.1 |
| DINO-v2-L [45] |
75.4 |
86.9 |
68.9 |
| DINO-v2-G [45] |
75.7 |
87.5 |
70.1 |
| SAM-B [25] |
59.8 |
76.9 |
60.1 |
| DINO-v1-B [6] |
53.5 |
76.9 |
54.2 |
| SD-2.1 [52] |
66.6 |
84.2 |
52.4 |
| CLIP-L [50] |
15.0 |
30.4 |
3.0 |
◼ means that we only obtain a trivial performance.
geometric |
visual |
meta |
COCO-20 |
FSS-1000 |
DAVIS-17 |
68.8 |
87.3 |
65.8 |
71.0 |
85.1 |
67.6 |
73.7 |
63.7 |
21.6 |
48.9 |
58.1 |
21.1 |
74.6 | 87.4 | 70.1 |
reversion |
negative |
COCO-20 |
FSS-1000 |
DAVIS-17 |
70.0 |
86.1 |
66.3 |
72.9 |
86.9 |
65.0 |
73.1 |
86.5 |
64.3 |
74.6 | 87.3 | 70.1 |
| semantic segmentation | instance segmentation | COCO-20 | FSS-1000 | DAVIS-17 | AVG | ||
COCO |
ADE20k |
COCO |
LVIS |
76.3 |
80.9 |
45.1 |
67.4 |
75.6 |
82.3 |
40.1 |
66.0 |
75.7 |
83.5 |
70.1 |
76.4 |
74.6 |
87.4 |
68.4 |
76.8 |
Generic semantic segmentation. We also evaluate SegIC on generic semantic segmentation. We use two widely-used semantic segmentation datasets, COCO [32] and ADE20k [67], for evaluation. To meet our in-context learning framework, we initially gather in-context examples from the training set along with the classes within each image and then perform segmentation in an in-context manner. As illustrated in Tab. 1, compared to specialist and generalist models for semantic segmentation, SegIC demonstrates strong performance under the settings that the classes are known in advance (marked with #). As shown in Tab. 1, our best model achieves 74.0 and 59.0 mIoU on COCO and ADE20k, respectively, surpassing the previous specialist and generalist models.
Open-vocabulary semantic segmentation. Similar to generic semantic segmentation, we can also extend SegIC to open-vocabulary semantic segmentation with the assistance of diffusion models. We employ in-context examples like one-shot segmentation to provide in-context examples for novel categories. Furthermore, we could also utilize StableDiffusion [52] to synthesize in-context examples like [44]. Compared to X-decoder [69], our method obtains competitive results on PC-459 and even outperforms previous methods on ADE-847.

### v3

# 2Shanghai Collaborative Innovation Center of Intelligent Visual Computing

@@ Abstract
In-context segmentation aims at segmenting novel images using a few labeled example images, termed as “in-context examples”, exploring content similarities between examples and the target. The resulting models can be generalized seamlessly to novel segmentation tasks, significantly reducing the labeling and training costs compared with conventional pipelines. However, in-context segmentation is more challenging than classic ones requiring the model to learn segmentation rules conditioned on a few samples. Unlike previous work with ad-hoc or non-end-to-end designs, we propose SegIC, an end-to-end segment-in-context framework built upon a single vision foundation model (VFM). In particular, SegIC leverages the emergent correspondence within VFM to capture dense relationships between target images and in-context samples. As such, information from in-context samples is then extracted into three types of instructions, i.e. geometric, visual, and meta instructions, serving as explicit conditions for the final mask prediction. SegIC is a straightforward yet effective approach that yields state-of-the-art performance on one-shot segmentation benchmarks. Notably, SegIC can be easily generalized to diverse tasks, including video object segmentation and open-vocabulary segmentation. Code will be available at https://github.com/MengLcool/SEGIC.
Keywords:
In-context learning Segmentation generalist Vision foundation model1 Introduction
@@
We design geometric, visual, and meta instructions that explicitly transfer knowledge from in-context samples to the target to facilitate in-context segmentation without tuning the parameters of vision foundation models.
SegIC demonstrates state-of-the-art performance on COCO-20i, FSS-1000 and recent LVIS-92i. Moreover, we conduct a comprehensive study on vision foundation models of various pre-text tasks, model sizes, and pre-training data. Furthermore, SegIC achieves competitive performance on novel tasks including video object segmentation and open-vocabulary segmentation, without ever seeing their training data.
@@ 2 Related Work
Vision foundation models. Recent years have witnessed great progress in large-scale vision pre-training [52, 10, 18, 8, 3, 17, 27], serving as the cornerstone for high-capacity foundation models. These pre-training approaches can be broadly categorized into two directions: vision-only pretraining [17, 18, 10, 8, 49, 3] and vision-language pre-training [52, 24, 33, 69, 56]. For vision-only pre-training, models aim to distinguish image/patch-level entities from different views [8, 10, 18] or reconstruct masked images [3, 18] from raw images. In vision-language pre-training, models strive to align cross-modal features into a unified visual-semantic space [52, 24, 69, 33], showcasing great open-set performance due to the transferable capabilities of language. Unlike these approaches that perform pre-training in an unsupervised or weakly supervised manner, SAM [27] is pre-trained on a huge amount of labeled segmentation data with precise prompts about locations. In this paper, we conduct extensive experiments using three types of pre-trained models as backbones to explore their potential for in-context segmentation. Interestingly, we observe that models with higher zero-shot semantic and geometric correspondence performance are more likely to be effectively utilized in our SegIC framework for in-context segmentation.
Unified vision downstream models. Unified vision downstream models have recently drawn significant attention due to their generalization capabilities and flexibilities. Unlike previous specialized vision models designed for specific datasets [55, 40, 19, 9, 64], vision generalists are tailored to handle multiple datasets [73, 43, 15, 44] and a wide range of tasks [74, 75, 23] within a single yet unified model. Recently, many studies [74, 75, 23, 27, 62, 63] have focused on developing techniques that unify segmentation tasks. In a similar spirit, our work also builds upon a unified output space for segmentation tasks. However, our goal is different—we aim to perform in-context segmentation that allows a model to effectively segment novel entities conditioned on a few samples.
@@
In-context learning. For the first time, GPT-3 [6] introduces a new learning paradigm known as in-context learning, which unifies various NLP tasks as text completion or question-answering tasks using provided prompts and examples. This approach enables language models to handle various tasks, including novel ones, by leveraging task examples, without requiring re-training or fine-tuning. Recent studies [62, 63, 38, 4, 1] explore this mechanism in vision tasks. Painter [62] and SegGPT [63] aim to achieve in-context segmentation through in-painting. They build upon the Mask Image Modeling (MIM) framework [17] to concatenate images and predictions into a 22 mosaic and make predictions by recovering the masked areas. In their pipelines, the vision backbone serves as both an image encoder and a mask decoder, which incurs significant computational costs. Moreover, they struggle to effectively leverage pre-trained models due to input shifts, leading to increased convergence challenges. Other approaches [71, 38] attempt using in-context segmentation via prompting SAM [27]. They build upon cross-image correspondences between in-context examples and target images by additional pre-trained models to generate prompts for SAM. However, these methods employ a two-stage pipeline, introducing redundancy and repeated computations. Consequently, if the model encounters limitations in one stage, it negatively impacts the final performance. In this work, we build an end-to-end in-context segmentation framework, leveraging the emergent correspondence using a single vision foundation model.
@@ 3.2 In-context Instruction Extraction
Visual instructions indicate the visual clues of the target entity. We utilize mask pooling [66] on reference image features to extract visual salient clues, :
By doing so, only information relevant to the reference mask, including low-level (texture, appearance) and high-level (semantic) clues, is used.
@@ 3.3 Mask Decoding
where indicate the projection layers for , respectively. , , and are the projected in-context instructions.
@@ 4.1 Training Data
COCO [35] is a widely used instance/semantic segmentation with 83K training samples and 41K samples of 80 categories. We use the 2014 split version to be consistent with the COCO-20i one-shot semantic segmentation benchmark.
@@ 4.2 Training Details
We implement SegIC in PyTorch and use 8 V100 GPUs for most of our experiments. We use a batch size of 32 (4 per GPU) in total, 1 point for point instruction, and 12544 points per mask to calculate mask loss following [11]. We merge the COCO instances for its semantic segmentation. Please refer to our supplementary material for the detailed training pipeline. Thanks to the frozen backbone, our method is extremely memory-efficient (with less than 10G memory cost in most of our experiments). SegIC is a single unified model that is jointly trained on mixed datasets, while evaluated on various datasets, separately. We utilize DINOv2 [49] as the default vision foundation model, with a ViT-B for all ablations, and ViT-L/G for the main experiments. We employ an AdamW [41] optimizer with a weight decay of We set an initial learning rate as and multiply 0.1 at the 10 epoch during training. Each dataset is sampled uniformly with 160K samples per epoch in the main experiments and 80K samples for the ablations. We perform data augmentations on target images and reference images respectively. We use large-scale jittering augmentation for semantic segmentation datasets, and normal data augmentations, including random resizing cropping, color jittering, and random horizontal flipping, for instance segmentation datasets. We random sample 1 mask per image for semantic segmentation datasets during training, while up to 10 masks for instance segmentation. The size of a single image is cropped/padded to 896896.
Method | Paraml | one-shot segmentation | video object segmentation | semantic seg | open-vocab seg | ||||||
COCO-20i | FSS-1000 | LVIS-92i | DAVIS-17 | YVOS-18 | COCO | ADE20k | PC-459 | A-847 | |||
mean mIoU | mIoU | mean mIoU | & | G | mIoU | mIoU | mIoU | mIoU | |||
few-shot seg specialist | |||||||||||
HSNet (RN50) [45] | 28M | 41.7 | 86.5 | 17.4 | n/a | n/a | n/a | n/a | n/a | n/a | |
VAT (RN50) [20] | 52M | 42.9 | 90.3 | 18.5 | n/a | n/a | n/a | n/a | n/a | n/a | |
FPTrans (B) [70] | 101M | 56.5 | - | - | n/a | n/a | n/a | n/a | n/a | n/a | |
VOS specialist | |||||||||||
AGAME (RN101) [25] | - | n/a | n/a | n/a | 70.0 | 66.0 | n/a | n/a | n/a | n/a | |
SWEM (RN50) [36] | 58M | n/a | n/a | n/a | 84.3 | 82.8 | n/a | n/a | n/a | n/a | |
XMem (RN50) [13] | 62M | n/a | n/a | n/a | 87.7 | 86.1 | n/a | n/a | n/a | n/a | |
semantic seg specialist | |||||||||||
MaskFormer (L) [12] | 212M | n/a | n/a | n/a | n/a | n/a | 64.8 | 54.1 | n/a | n/a | |
Mask2Former (L) [11] | 216M | n/a | n/a | n/a | n/a | n/a | 67.4 | 56.1 | n/a | n/a | |
MaskDINO (L) [31] | 225M | n/a | n/a | n/a | n/a | n/a | - | 56.6 | n/a | n/a | |
segmentation generalist | |||||||||||
OneFormer (L) [23] | 237M | n/a | n/a | n/a | n/a | n/a | 67.4 | 57.7 | - | - | |
UNINEXT (L) [68] | 340M | n/a | n/a | n/a | 77.2 | 78.1 | - | - | - | - | |
X-decoder (L) [74] | 341M | n/a | n/a | n/a | n/a | n/a | 67.5 | 58.1 | 29.6 | 9.2 | |
SEEM (L) [75] | 341M | - | - | - | 58.9 | 50.0 | 67.6 | - | - | - | |
in-context generalist | |||||||||||
Painter (L) [62] | 354M | 33.1 | 61.7 | 10.5 | 34.6 | 24.1 | - | 49.9 | - | - | |
SegGPT (L) [63] | 354M | 56.1 | 85.6 | 18.6 | 75.6 | 74.7 | - | *39.6 | - | - | |
PerSAM† (H) [71] | 0 | 23.0 | 71.2 | 11.5 | 60.3 | - | n/a | n/a | n/a | n/a | |
PerSAM-F† (H) [27] | 2 | 23.5 | 75.6 | 12.3 | 71.9 | - | n/a | n/a | n/a | n/a | |
Matcher† (H+G) [38] | 0 | 52.7 | 87.0 | 33.0 | 79.5 | - | n/a | n/a | n/a | n/a | |
SegIC (L) | 5M | 76.1 | 86.8 | 44.6 | 71.4 | 62.7 | #72.9 | #55.5 | #33.5 | #18.9 | |
SegIC (G) | 5M | 74.5 | 88.4 | 47.8 | 73.7 | 65.4 | #74.0 | #59.0 | #34.9 | #20.1 |
@@ 4.3 Main Results
For the main experiments, we compare SegIC with other specialist/generalist methods on several benchmarks of different segmentation tasks. We use SegIC of DINOv2 [49] of large and giant versions as the backbone for the main experiments.
One-shot semantic segmentation. To demonstrate the generalization capability from known categories to unknown ones, we evaluate SegIC on one-shot semantic segmentation benchmarks. Following SegGPT [63], we evaluate SegIC in two one-shot semantic segmentation settings: in-domain using COCO-20i[47] (the training set of COCO-20i is a subset of ours) and out-of-domain with FSS-1000 [34] (without seeing any training samples of FSS-1000). As depicted in Tab. 1, SegIC has achieved state-of-the-art performance in both of these settings. To ensure a fair comparison, we report in-domain performance of specialist models for COCO-20i reported in [63]. Additionally, for FSS-1000, we highlight the performance trained on FSS-1000 in gray and zero-shot results in black. Notably, SegIC outperforms previous generalist models by a significant margin (more than 20 of mean mIoU) on COCO-20i and achieves competitive results that are very close to specialist models on FSS-1000, even without ever being trained on it. Furthermore, we conduct experiments on LVIS-92i [38], a more challenging one-shot benchmark built upon LVIS [16]. On LVIS-92i, SegIC surpasses Matcher, the previous SoTA, by a large margin (from 33.0 to 47.8).
Method | F0 | F1 | F2 | F3 | mean |
HSNet [45] | 37.2 | 44.1 | 42.4 | 41.3 | 41.2 |
VAT [20] | 39.0 | 43.8 | 42.6 | 39.7 | 41.3 |
FPTrans [70] | 44.4 | 48.9 | 50.6 | 44.0 | 47.0 |
MSANet [22] | 47.8 | 57.4 | 48.7 | 50.5 | 51.1 |
SegIC | 55.8 | 54.7 | 52.4 | 51.4 | 53.6 |
@@
Since we include the evaluation categories in the training process for COCO-20i in Tab. 1, for a rigorous comparison, we follow its training setting that trains and tests our model on 4 splits [47] separately, avoiding seeing categories for evaluation. As shown in Tab. 2, our method still achieves state-of-the-art performance on COCO-20i. Furthermore, as shown in the last row of Tab. 2, by joint training on COCO-excluded datasets (including ADE20k, LVIS, and FSS-1000), there is a significant performance gain across all splits. This further demonstrates the effectiveness of our in-context generalization capabilities.
@@
Zero-shot video object segmentation. Video object segmentation (VOS) aims to segment specific objects in video frames. In this work, we focus on the semi-supervised VOS setting [51, 67], where the masks that appeared first time are given as references. We evaluate SegIC without any fine-tuning on video datasets to demonstrate our generalization capabilities. We choose two commonly used VOS datasets: DAVIS-17 [51] and YouTube-VOS-18 [67]. We export two metrics commonly used in VOS for evaluation: the score for DAVIS-17 and score for YouTube-VOS-18, with their official evaluation servers or toolkits. As shown in Tab. 1, when compared to VOS specialist models, SegIC achieves competitive performance on VOS benchmarks, even without seeing any training videos. Furthermore, in comparison to segmentation generalist models, SegIC surpasses Painter [62], SEEM [75], and PerSAM [71] by a significant margin, and competes favorably with recent generalist models [63, 38]. Additionally, the VOS task pipeline in SegIC is relatively simple. We do not use any test time augmentation tricks (TTA) used in [13]. Moreover, it does not involve dense feature interaction at the patch level, as in SegGPT, and does not require the use of a pre-trained SAM for segmentation.
Generic semantic segmentation. We also evaluate SegIC on generic semantic segmentation benchmarks, which need to segment dataset-dependent pre-defined categories within each image. We adapt generic semantic segmentation to our in-context learning framework by gathering in-context examples from the training set and then performing segmentation in an in-context manner. We use two widely-used semantic segmentation datasets, COCO [35] and ADE20k [72], for evaluation. As illustrated in Tab. 1, compared to specialist and generalist models for semantic segmentation, SegIC demonstrates strong performance under the settings that the classes are known in advance (marked with #). As shown in Tab. 1, our best model achieves 74.0 and 59.0 mIoU on COCO and ADE20k, respectively, surpassing the previous specialist/generalist models.
Open-vocabulary semantic segmentation. Similar to generic semantic segmentation, we can also extend SegIC to open-vocabulary semantic segmentation. Furthermore, we could also utilize StableDiffusion [56] to synthesize images with coarse masks for those categories lacking in-context examples like [48]. Compared to X-decoder [74], our method obtains competitive and even better results on PC-459 and ADE-847.
Encoder | Arch | HRes | MRes | Taskp | Datap | COCO-20i | FSS-1000 | DAVIS-17 | AVG |
DINOv2-B [49] | ViT | ✓ | ✓ | image/patch self-distillation | 142M | 74.6 | 87.4 | 68.4 | 76.8 |
DINOv2-L [49] | ViT | ✓ | ✓ | image/patch self-distillation | 142M | 75.4 | 86.9 | 68.9 | 77.1 |
DINOv2-G [49] | ViT | ✓ | ✓ | image/patch self-distillation | 142M | 75.7 | 87.5 | 70.1 | 77.8 |
DINOv1-B [8] | ViT | ✓ | image self-distillation | 1M | 53.5 | 76.9 | 54.2 | 61.5 | |
SAM-B [27] | ViT | ✓ | interactive segmentation | 11M | 59.8 | 76.9 | 60.1 | 65.6 | |
SD-2.1 [56] | UNet | ✓ | ✓ | text-to-image generation | 2B | 66.6 | 84.2 | 52.4 | 67.7 |
OpenCLIP-ConvNext-B [21] | CNN | ✓ | image-text alignment | 400M | 65.1 | 77.0 | 50.4 | 64.2 | |
CLIP-ViT-B [52] | ViT | image-text alignment | 400M | 15.3 | 29.4 | 1.3 | 15.3 | ||
MAE-B [17] | ViT | mask image modeling | 1M | 12.0 | 23.7 | 2.3 | 12.6 | ||
ConvNext-B [39] | CNN | ✓ | image classification | 1M | 62.0 | 72.8 | 37.6 | 57.5 |
4.4 Rethinking Vision Foundation Model Pre-training
Existing works improve VFMs by scaling up data and model sizes, modifying the model architectures, and utilizing different pre-text tasks. It is unknown how much those factors affect the performance when the encoder is frozen. We delve deeper into the potential of various frozen pre-trained VFMs [8, 49, 17, 27, 56, 52, 21, 39]. For each VFM, we freeze it to extract image features and dense correspondences then map them into the hidden space for mask decoding, equally. Additionally, we assess the zero-shot semantic correspondence score to underscore the emerging downstream capabilities from different perspectives directly. More specifically, we evaluate a percent of correct keypoints score (PCK) on SPair-71k [46] following to [14] with cosine similarity between the dense feature maps of the two images for semantic correspondence. We observe that (1) high/multi-resolution is essential; (2) pre-text tasks are important; (3) model and data scales do not necessarily help for ICL segmentation; (4) zero-shot correspondence can imply the ICL segmentation capacity.
High/multi-resolution is essential. It is evident in Tab. 3 that the models (CLIP-ViT-B and MAE-B), lacking support for high-resolution or multi-resolution only exhibit trivial performance within our frozen-backbone framework. Despite the remarkable success of vision transformers, they encounter limitations in directly adapting to various input resolutions without fine-tuning due to fixed-length visual tokens and position embeddings. Unlike vision transformers, CNNs, with their stacked convolution design, demonstrate seamless adaptation of parameters to images of different resolutions, even only pre-trained on low-resolution images. To alleviate this limitation, DINO-v1/v2 enhance vanilla ViTs with multi-resolution inputs and position embedding interpolation during pre-training. Consequently, they achieve superior segmentation performance. Overall, the observations indicate that the high/multi-resolution capabilities of the pre-trained VFMs play a key role in in-context segmentation.
Pre-text tasks are important. Once the conditions for supporting high-resolution inputs are met, pre-text tasks employed during pre-training and the scale of data also impact the in-context segmentation capacities. Illustrated in Fig. 2, with the same model size and pre-training data, DINOv1 outperforms those classification-pre-trained models in both zero-shot semantic correspondence and in-context segmentation tasks by a large margin. This emphasizes the substantial influence of the pre-text task on downstream capabilities.
Model and data scales do not directly aid ICL segmentation. Equipping with larger-scale data and advanced self-distillation techniques, DINOv2 outperforms DINOv1 by a considerable margin. However, as shown in Tab. 3, scaling up the model size from base (B) to giant (G), is not a clear performance boost. Furthermore, compared to those pre-trained on than 10 times data, e.g. OpenCLIP-ConvNext-B and SD-2.1, DINOv2 still exhibits significant advantages. Those demonstrate that model and data scales do not directly help.
Zero-shot correspondence implies the segmentation capacity. We further study the relationship between correspondence and segmentation performance with different backbones. As shown in Fig. 2, the performance on segmentation is proportional to correspondence, demonstrating a strong consistency between these two tasks. It further confirms our motivation to leverage the emergent correspondence for in-context segmentation. This insight further inspires us, suggesting that pre-training emphasizing inter/intra-image correspondence may lead to better in-context segmentation potentials, e.g. the image/patch-level discriminative self-distillation [49].